Make a study about how many sheets of paper you consumed weekly in 

answering your Self Learning Modules. Record the quantity (total number of sheets)

per subject, then construct a probability distribution. Compute the mean, variance, 

and the standard deviation of the probability distribution you made. Interpret the 

result, then find out how many weeks you will consume 50 sheets of pad paper


1
Expert's answer
2021-04-14T00:12:01-0400

Math - 10 sheets

Physics -8

Literature - 15

Biology -7

Chemistry - 8

Computer Science - 11

English -12

History - 10


Total number of sheets:

"N=10+8+15+7+8+11+12+10=81"

Number of subjects:

"n=8"

Probability:

"P(Math)=10\/81"

"P(Physics)=8\/81"

"P(Literature)=5\/27"

"P(Biology)=7\/81"

"P(Chemistry)=8\/81"

"P(Computer Science)=11\/81"

"P(English)=12\/81=4\/27"

"P(History)=10\/81"


The mean:

"\mu=81/8=10.125"

Variance:

"V=\frac{\sum(x_i-\mu)^2}{N}"

"V=\frac{0.125^2+2.125^2+4.875^2+3.125^2+2.125^2+0.875^2+1.875^2+0.125^2}{81}=0.579"

Standard deviation:

"\sigma=\sqrt{0.579}=0.76"


The mean is average number of sheets per subject. Variance and standard deviation define dispersion of variable (number of sheets per subject) respect to the mean.


I will consume 50 sheets of pad paper in

"50/81=0.62\ week"